Per Capita GDP Economy A Economy B GDP:             400000   442200 GDP:              100000  &nb

Per Capita GDP

Economy A

Economy B

GDP:             400000   442200

GDP:              100000    124000

Population:   8              8.04                    

Population: 20            24

1.Which economy grew?

2.Which economy grew more in $?

3.Which economy grew more in %?

4.Which has more people?

5.Which population grew?

6.Which population grew more in percent?

7.Which added more people?

8.Which has more per capita?

9.Which grew in per capita terms?

10.Which grew more in $ per capita?

11.Which grew more in % per capita?

12.Which is happier?
Answer #1

1. As the GDP of both nation A and B rose hence both economies grew in size.

2. Change in GDP for nation A: 442200 - 400000 = 42200

  Change in GDP for nation B: 124000- 100000= 24000

Change in GDP for nation A is greater then Change in GDP for nation B

3. % Change in GDP for nation A; 42200/400000 = 10.55 %

% Change in GDP for nation B; 24000/100000 = 24 %

Hence % Change in GDP for nation B is greater then % Change in GDP for nation A